a sum of money amounting to 5.55 consistsof 25cents and 1 peso coins.If there are 27 coins in all, how many each kind are there.

Answer by CubeyThePenguin(3113)

p = number of 1-peso coins
q = number of 25-peso coins

p + q = 27
p + 25q = 555

Multiply the first equation by 25. We are planning to eliminate the variable q.

25p + 25q = 675
p + 25q = 555

Subtract to get 24p = 120, so p = 5.

There are 5 pennies and 22 quarters.